Connie can tpe a report in 4 hours. Mika can type the same report in two hours. How long would it take the two of them, working together, to type the report?
Answer by ReadingBoosters(3246)

Rate(Connie) 1 report / 4 hours
Rate(Mika) 1 report / 2 hours
Rate(Connie & Mika) = 1/4 + 1/2 = 1/x where x is their rate together
...
1/4 + 1/2 = 1/x
1/4 + 1/2(2/2) = 1/x
1/4 + 2/4 = 1/x
3/4 = 1/x
Knowing 1/x is the inverse of x, invert the other side also
4/3 = x = 1 1/3 hours = 1.33 hours
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com